OpenStudy (anonymous):

Simplify the trigonometric expression 1/1+sin0 + 1/1-sin0

OpenStudy (agent0smith):

Use parentheses :P the denominator needs them

OpenStudy (anonymous):

\[\frac{ 1 }{ 1+\sin \theta }+ \frac{ 1 }{ 1-\sin }\]

OpenStudy (agent0smith):

Wow good job.

OpenStudy (agent0smith):

Get common denominators:\[\large \frac{ 1(1-\sin \theta) }{ (1+\sin \theta) (1-\sin \theta) }+ \frac{ 1 (1+\sin \theta)}{ (1-\sin \theta) (1+\sin \theta) }\]

OpenStudy (anonymous):

\[\frac{ 1}{ 1+\sin \theta }+\frac{ 1}{ 1-\sin \theta }\]

OpenStudy (anonymous):

thats what i meant ^

OpenStudy (anonymous):

the answer is 2 right?

OpenStudy (imstuck):

you need a common denominator here. Multiply the left expression by (1-sintheta/1-sintheta), and multiply the expression on the right by (1+sin theta/1+ sin theta), like agent0smith did.

OpenStudy (agent0smith):

\[\large \frac{ 1(1-\sin \theta) }{ (1+\sin \theta) (1-\sin \theta) }+ \frac{ 1 (1+\sin \theta)}{ (1-\sin \theta) (1+\sin \theta) } =\]\[\large \frac{2}{ (1-\sin \theta) (1+\sin \theta) } = \frac{2}{ 1-\sin^2 \theta}\]

OpenStudy (imstuck):

There is an identity hidden in the 1 - sin^2 theta...

OpenStudy (imstuck):

1 - sin^2 theta = cos^2 theta.

OpenStudy (imstuck):

\[\frac{ 2 }{ \cos ^{2} }\]

OpenStudy (agent0smith):

Oh yeah \[\large \frac{2}{ 1-\sin^2 \theta} = \frac{ 2 }{ \cos^2 \theta }= 2 \sec ^2 \theta\]

OpenStudy (imstuck):

Cool! This just goes on and on.... ; )

OpenStudy (anonymous):

thank you